# Murphy: Multi-Turn GRPO for Self Correcting Code Generation

### v1

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) has emerged as a powerful framework for enhancing the reasoning capabilities of large language models (LLMs). However, existing approaches such as Group Relative Policy Optimization (GRPO) and its variants, while effective on reasoning benchmarks, struggle with agentic tasks that require iterative decision-making. We introduce Murphy, a multi-turn reflective optimization framework that extends GRPO by incorporating iterative self-correction during training. By leveraging both quantitative and qualitative execution feedback, Murphy enables models to progressively refine their reasoning across multiple turns. Evaluations on code generation benchmarks with model families such as Qwen and OLMo show that Murphy consistently improves performance, achieving up to a $8\%$ relative gain in pass@1 over GRPO, on similar compute budgets.
@@ 1 Introduction
Qwen3-1.7BQwen3-4BOLMo-2-7B-Ins02244668810102.82.82.12.13.23.26.66.62.92.97.47.4ModelsPercentage (%)($\%$)GRPOMurphy
Figure 1: Percentage change in coding problems solved by models trained with Murphy and GRPO over the base model across three models and datasets. Murphy-trained models solve up to $4.2\%$ more problems than GRPO. See Tab. 1 for details.
A growing body of work explores large language models (LLMs) as software engineering agents that interact with their environment through code execution and feedback (Shinn et al., 2023; Lingam et al., 2025; Miserendino et al., 2025). Rather than producing a single static response, these systems operate within agentic scaffolds (Yao et al., 2023; Zhou et al., 2024) that guide iterative reasoning and allow LLMs to act, observe, and improve over multiple rounds of interaction. For example, in a typical coding agnetic scaffold (Shinn et al., 2023; Lingam et al., 2025), the agent generates a solution by performing a single or series of actions, and executes it for evaluation, often through unit tests or other automated checks. When execution fails, the agent receives feedback such as error messages, stack traces, or failing test cases, and is re-prompted with the original task, its previous attempt, and the new feedback. This process continues for multiple turns until the model succeeds or reaches a fixed iteration limit. These systems highlight the growing ability of LLMs to reason and adapt through environmental feedback at inference time. In code generation tasks (Jiang et al., 2025), such feedback naturally arises from execution logs, compiler errors, or test results. However, these methods (Shinn et al., 2023; Lingam et al., 2025) remain fundamentally training-free: they improve model behavior through structured inference via feedback rather than through parameter updates.
@@
Main Contributions. 1. We introduce Murphy, a multi-turn RLVR algorithm that extends GRPO to leverage execution feedback for grounded reasoning and self-correction in code generation tasks. (Sec. 4) 2. We design pruning mechanisms that enable scalable multi-turn RLVR training (Murphy) while preserving performance. (Subsec. 4.1) 3. We evaluate Murphy on three code generation benchmarks across two model families (OLMo, Qwen) and sizes (1.7B–7B), achieving up to $5\%$ pass@1 over GRPO baselines. (Sec. 5)
@@ 3 Background: GRPO
Group Relative Policy Optimization (GRPO; (Shao et al., 2024)) is a variant of Proximal Policy Optimization (PPO) designed to improve the efficiency and stability of policy updates in LLM fine-tuning. Unlike PPO, which estimates advantages using a learned value function (critic), GRPO replaces the critic with an empirical baseline: the mean reward across all generations produced by the model for the same prompt. In both methods, rewards are provided by a reward model. Specifically, for each input prompt, the model samples a set of $G$ candidate responses, forming a response group. The reward model assigns a score to each response, and advantages are computed by standardizing rewards within the group: subtracting the group mean and dividing by the group standard deviation. This yields relative, normalized advantage values. As in PPO, GRPO may incorporate a penalty term to prevent the updated policy from drifting too far from the reference policy, typically enforced via a Kullback–Leibler (KL) divergence regularizer to ensure stable updates. A formal definition of the GRPO objective, along with additional details, is provided in App. B.
@@ 4 Proposed Method: Murphy
Figure 2: Overview of Murphy. Given an input prompt (q), $G$ code generations (o) are generated and evaluated using a reward function (r). Generations that do not achieve the maximum reward are revised based on executor feedback (f), combining the original prompt with the failed output, and re-prompted to generate another $G$ candidates. This iterative process continues for a fixed number of turns, with rewards from later turns propagated backward. The example illustrates the case with $G=2$, where $G$ represents the number of rollouts per prompt, and $\rho(\cdot)$ denotes the credit assignment strategy.
Extending GRPO to multi-turn interaction requires formalizing how environmental feedback informs optimization. We study a feedback-rich code generation setting where model outputs are executed and scored against test suites, yielding two feedback types: (1) quantitative (e.g., proportion of test cases passed) and (2) qualitative (e.g., error traces or failing cases). Murphy extends GRPO by introducing (i) a multi-turn rollout mechanism that conditions generation on feedback, and (ii) a credit-assignment scheme that propagates rewards from later successful turns to earlier attempts.
Multi-turn rollout: For each task prompt, the policy begins at turn 1 by generating $G_{1}$ candidate solutions, each evaluated for reward against the prompt’s test suite. Failed generations are paired with their corresponding feedback and, together with the original prompt and prior outputs, form new conditioning contexts for subsequent turns. At turn $s$, the model generates $G_{s}$ new candidates for each failed case, continuing this feedback–refinement process for a fixed number of turns. This per-prompt, iterative rollout enables progressive improvement conditioned on feedback. Unlike standard GRPO, Murphy delays advantage computation until the full rollout is complete, allowing final rewards to retroactively shape credit assignment.
Credit assignment: Once rewards for all turns are computed, Murphy propagates future rewards backward to earlier turns using a temporal credit-assignment scheme. Although early generations may achieve low initial rewards, incorporating execution feedback in later turns often leads to successful refinements. To ensure that these intermediate steps are properly credited, rewards from later successful generations are distributed to earlier ones according to their contribution to eventual success. After reward redistribution, advantages are computed for each generation, and the Murphy loss is applied to update the policy. Together, these mechanisms extend GRPO to handle multi-turn, feedback-conditioned optimization. Below, we formalize the multi-turn rollout mechanism and credit-assignment process and define the Murphy objective, extending GRPO to handle multi-turn feedback.
@@
We denote the model policy by $\pi_{\theta}(\cdot\mid\cdot)$ and the reference (older) policy by $\pi_{\theta_{\text{old}}}(\cdot\mid\cdot)$. Let $\mathcal{P}(Q)$ the distribution over input prompts/questions $Q$, and $O$ the output space. As described earlier, in the first turn, the model generates $G_{1}$ candidate solutions for each prompt. For generations that fail, feedback is obtained from the environment. The model is then re-prompted with the original prompt, its previous output, and the corresponding feedback to produce $G_{s}$ new generations per prompt at turn $s$. This iterative procedure naturally forms a tree structure, where the root corresponds to the original prompt, and each subsequent generation (augmented with feedback) forms a child node, with generations at turn $s+1$ linked to their parent at turn $s$ (see Fig. 2). We formalize this multi-turn rollout process as follows.
@@
Multi-turn rollout formalism.We define a feedback-conditioned rollout tree that captures how model generations evolve across multiple turns of interaction with the environment. Let $s$ denote the turn index, $S$ the total number of turns, and $G_{s}$ the number of generations per prompt at turn $s$.
Turn 1:At the first turn ($s=1$), the model receives an input prompt $q_{(1)}$ sampled from $\mathcal{P}(Q)$ and generates $G_{1}$ candidate programs which forms a response group $\{o_{\{q_{(1)},1\}},o_{\{q_{(1)},2\}},\ldots,o_{\{q_{(1)},G_{1}\}}\}$. Note that $\{o_{\{q_{(1)},j\}}\}_{j=1}^{G_{1}}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q_{(1)})$ . For any generation indexed by $j$, the generated code is executed against its associated test suite, producing a numerical reward $r_{\{q_{(1)},j\}}$ and environment feedback $f_{\{q_{(1)},j\}}$. In our setting, the reward represents the proportion of test cases passed, while the environment feedback consists of the specific unit tests that passed or failed, along with any corresponding error messages. These generations together form the first layer of output nodes in the rollout tree.
Turn $s$:Building on the previous turn, this procedure extends recursively to any turn $s\in\{1,\dots,S-1\}$. Note that we define:
$i_{[1:s]}=i_{1},\ \dots,\ i_{s}$
For each generation $o_{\{q_{(s,i_{1:(s-1)})},i_{s}\}}$ that fails to achieve the maximum reward, the feedback-conditioned prompt is constructed:
$q_{(s+1,i_{[1:s]})}=[\,q_{(s,i_{[1:s-1]})},o_{\{q_{(s,i_{1:(s-1)})},i_{s}\}},f_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}]$
where $\cdot\,$] denotes textual concatenation. The model is then re-invoked to generate $G_{s+1}$ new candidate solutions $\{\,o_{\{q_{(s+1,i_{[1:s]})},k\}}\,\}_{k=1}^{G_{s+1}}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q_{(s+1,i_{[1:s]})})$. Each generation is evaluated to obtain $(r_{\{q_{(s+1,i_{[1:s]})},k\}},\,f_{\{q_{(s+1,i_{[1:s]})},k\}})$, which denote the corresponding reward and feedback, respectively. The resulting generations, $\{o_{\{q_{(s+1,i_{[1:s]})},k\}}\,\}_{k=1}^{G_{s+1}}$ form the children nodes of the unique corresponding parent node $o_{\{q_{(s,i_{1:(s-1)})},i_{s}\}}$. Hence, a complete path from the root (at the prompt of turn 1) to a leaf at stage $S$ (the output of the model), is expressed as:
$q_{(1)}\rightarrow o_{\{q_{(1)},i_{1}\}}\rightarrow q_{(2,i_{1})}\rightarrow o_{\{q_{(2,i_{1})},i_{2}\}}\rightarrow\cdots\rightarrow o_{\{q_{(S,i_{1},\dots,i_{S-1})},i_{S}\}}$
Note that leaf nodes at turn $S$ represent the final generations obtained after completing all refinement steps. Once rewards for all turns are computed, the rewards from these terminal nodes are propagated backward through their parent nodes according to the credit assignment strategies described below.
@@
Credit assignment formalism.After all outputs and corresponding rewards are generated and the tree structure is constructed, we focus on assigning credit from later turns back to earlier ones. To achieve this, we employ two distinct strategies.
@@
Max Reward Strategy (MaRS):The Max Reward Strategy is defined recursively, proceeding from the final turn back to the root. Since the final turn $S$ has no children, the rewards at this turn remain unchanged. We then consider turn s=S1s=S-1. Let $o_{\{q_{(s,i_{1:(s-1)})},i_{s}\}}$ denote a generation at turn s=S1s=S-1 with an associated reward $r_{\{q_{(s,i_{1:(s-1)})},i_{s}\}}$. If this generation already achieves the maximum reward, it has no children; otherwise, its children are defined as:
$C(o_{\{q_{(s,i_{1:(s-1)})},i_{s}})=\{o_{\{q_{(s+1,i_{[1:s]})},1\}},\ \ldots,\ o_{\{q_{(s+1,i_{[1:s]})},G_{S}\}}\}$
@@
$r_{\{q_{(s,i_{1:(s-1)})},i_{s}\}}=\text{max}\Bigg(r_{\{q_{(s,i_{1:(s-1)})},i_{s}\}},\ \text{max}\Big(C_{r}(o_{\{q_{(s,i_{1:(s-1)})},i_{s}})\Big)\Bigg)$
@@
This rule selects the greater value between a node’s own reward and the maximum reward among its descendants. Intuitively, the inner maximum represents the best improvement achievable through feedback, while the outer maximum guarantees that a node’s reward does not decrease when feedback fails to yield better performance. This formulation therefore captures the maximum attainable progress achievable from feedback at each turn. The procedure is applied recursively: rewards are first updated for all generations at turn S1S-1, then propagated backward toward the root. At each turn $s$, rewards are reassigned based on the updated values from turn $s+1$, continuing until all turns have been processed.
Mean Reward Strategy (MeRS):The Mean Reward Strategy follows the same recursive credit assignment structure as the Max Reward Strategy (MaRS), with the key difference lying in how rewards are propagated. Instead of using the maximum reward among descendants, MeRS updates each node’s reward using the mean of its child rewards, normalized over the number of stages the update has taken place over. Specifically, for a generation $o_{\{q_{(s,i_{1:(s-1)})},i_{s}\}}$ at stage $s$, the update rule is,
$r_{\{q_{(s,i_{1:(s-1)})},i_{s}\}}=\frac{1}{S-s+1}\Bigg(r_{\{q_{(s,i_{1:(s-1)})},i_{s}\}}+\gamma\cdot\text{mean}\Big(C_{r}(o_{\{q_{(s,i_{1:(s-1)})}\}}\Big)\Bigg)$
where $\gamma$ is a discount factor controlling the influence of descendant rewards, and the normalization term $\frac{1}{S-s+1}$ ensures that credit assignment remains consistent across varying depths of the tree. Intuitively, this formulation captures the average improvement achieved through feedback, scaled by how many stages of updates have occurred. The normalization prevents early-stage nodes (with many descendants) from accumulating disproportionately large credit, while ensuring that terminal nodes, having no children, are not penalized. All other aspects of the recursive procedure, including the bottom-up updates from stage S1S-1 to the root, remain identical to those in MaRS.
@@
MaRS vs MeRS:MaRS propagates the maximum descendant reward to the earlier stages, emphasizing peak performance, whereas MeRS propagates the mean reward, emphasizing stability and overall consistency. MaRS captures best-case improvement, while MeRS provides a smoother estimate of expected progress. Together, they offer complementary views of feedback-driven credit assignment.
@@
Murphy Objective:Once the rewards are reassigned according to the chosen credit assignment strategy (MaRS or MeRS), advantages are computed similar to standard GRPO. The adjusted rewards serve as the basis for computing normalized advantages at each turn. Conditioned on a prompt $\tilde{q}$ and for $G_{s}$ generations, we normalize each reward by subtracting the mean reward and dividing by the standard deviation of the rewards obtained across the $G_{s}$ generations, yielding the normalized advantage $\hat{A}^{\textsc{Murphy}}_{\tilde{q},i,t}$. Additionally, as defined earlier, each $i$-th generation at stage $s$, $o_{\{q_{(s,i_{1:(s-1)})},i_{\}}}\in O$, corresponds to a complete output trajectory, where $o_{\{q_{(s,i_{1:(s-1)})},i_{,}t\}}$ denotes the $t$-th token and $o_{\{q_{(s,i_{1:(s-1)})},i,<t\}}$ the prefix up to (but excluding) token $t$. We denote the sequence length of the $i$-th generation by $|o_{\{q_{(s,i_{1:(s-1)})},i\}}|$. At each turn, the GRPO objective is applied using these credit-adjusted advantages. This per-turn optimization allows Murphy to incorporate feedback from later turns into earlier updates, effectively extending GRPO to a multi-turn setting. Finally, the divergence between the current and reference policies is captured by $D_{\mathrm{KL}}(\pi_{\theta}\,\|\,\pi_{{\text{ref}}})$, computed over all tokens in the generated sequences. The resulting optimization objective, which integrates credit-assigned rewards, normalized advantages, and KL regularization at each turn, defines the Murphy objective, distinguishing it from the standard GRPO formulation. The full Murphy objective is presented in Sec. 4.
@@
Definition 1. (Murphy Objective) $\mathcal{J}_{\textsc{Murphy}}(\theta)$ $=\mathbb{E}_{q\sim\mathcal{P}(Q)}[\sum_{s,i_{1},\ldots i_{S}}\mathcal{J}_{q_{(s,i_{1:s})}}(\theta)]$ Where the per prompt objective at each turn $s$ is: $\mathcal{J}_{q_{(s,i_{1:(s-1)})}}(\theta)=\$$\mathbb{E}_{\{o_{(\tilde{q},i)}\}_{i=1}^{G_{s}}\sim\pi_{\theta_{\text{old}}}(O|\tilde{q})}\Bigg[\sum_{i=1}^{G_{s}}\frac{1}{G_{s}|o_{\{\tilde{q},i\}}|}\sum_{t=1}^{|o_{\{\tilde{q},i\}}|}\Bigg(\min\Big(R_{\theta}(\tilde{q},i,t)\hat{A}^{\textsc{Murphy}}_{\tilde{q},i,t},$ $\quad\text{clip}\left(R_{\theta}(\tilde{q},i,t),\,1-\epsilon,\,1+\epsilon\right)\hat{A}^{\textsc{Murphy}}_{\tilde{q},i,t}\Big)-\beta D_{\mathrm{KL}}(\pi_{\theta}\,\|\,\pi_{{\text{ref}}})\Bigg)\Bigg]$ with, $\tilde{q}$ $=q_{(s,i_{1:(s-1)})}$ $R_{\theta}(\tilde{q},i,t)$ $=\frac{\pi_{\theta}(o_{\{\tilde{q},i,t\}}\mid\tilde{q},o_{\{\tilde{q},i,<t\}})}{\pi_{\theta_{\text{old}}}(o_{\{\tilde{q},i,t\}}\mid\tilde{q},o_{\{\tilde{q},i,<t\}})}$ And, $\hat{A}^{\textsc{Murphy}}_{\tilde{q},i,t}$ denotes the advantage.

### v2

# Murphy: Multi-Turn GRPO for Self-Correcting Code Generation

@@ Abstract
Reinforcement Learning with Verifiable Rewards (RLVR) has emerged as a powerful framework for enhancing the reasoning capabilities of large language models (LLMs). However, existing approaches such as Group Relative Policy Optimization (GRPO) and its variants, while effective on reasoning benchmarks, struggle with agentic tasks that require iterative decision-making. We introduce Murphy, a multi-turn RLVR framework that incorporates execution feedback directly into training, extending GRPO to optimize over multi-turn trajectories where models iteratively refine solutions. Murphy combines a feedback-conditioned rollout tree with trajectory-level credit assignment, and uses pruning to reduce the cost of multi-turn optimization. Evaluations on code generation benchmarks with two model families show that Murphy consistently improves multi-iteration performance, achieving up to an $8\%$ absolute gain in pass@1 over compute-matched GRPO baselines, and outperforming the prior leading method that incorporates multi-turn execution feedback.
Machine Learning, ICML
@@ 1 Introduction
Qwen3-1.7BQwen3-4BOLMo-2-7B-Ins011223344556677882.82.81.41.43.23.26.76.72.92.97.47.4ModelsPercentage (%)($\%$)GRPOMurphy
Figure 1: Percentage change in coding problems solved by models trained with Murphy and GRPO over the base model across three models and datasets. For Qwen3-4B, we additionally include results on the Aider Polyglot benchmark. Murphy-trained models solve up to $4.2\%$ more problems than GRPO. See Tab. 1 and Subsec. 5.1 for details.
A growing body of work explores large language models (LLMs) as software engineering agents that interact with their environment through code execution and feedback (Shinn et al., 2023; Lingam et al., 2025; Miserendino et al., 2025). Rather than producing a single static response, these systems operate within agentic scaffolds (Yao et al., 2023; Zhou et al., 2024) that guide iterative reasoning and allow LLMs to act, observe, and improve over multiple rounds of interaction. For example, in a typical coding agentic scaffold (Shinn et al., 2023; Lingam et al., 2025), the agent generates a solution by performing a single or series of actions, and executes it for evaluation, often through unit tests or other automated checks. When execution fails, the agent receives feedback such as error messages, stack traces, or failing test cases, and is re-prompted with the original task, its previous attempt, and the new feedback. This process continues for multiple turns until the model succeeds or reaches a fixed iteration limit. These systems highlight the growing ability of LLMs to reason and adapt through environmental feedback at inference time. In code generation tasks (Jiang et al., 2025), such feedback naturally arises from execution logs, compiler errors, or test results. However, these methods (Shinn et al., 2023; Lingam et al., 2025) remain fundamentally training-free: they improve model behavior through structured inference via textual feedback rather than through parameter updates.
@@
[…]
This work focuses on code generation tasks where execution provides structured, verifiable feedback, enabling us to measure feedback-driven refinement using agentic scaffolds at evaluation time.
@@ 3 Background: GRPO
Group Relative Policy Optimization (GRPO; (Shao et al., 2024)) is a variant of Proximal Policy Optimization (PPO) designed to improve the efficiency and stability of policy updates in LLM fine-tuning. Unlike PPO, which estimates advantages using a learned value function (critic), GRPO replaces the critic with an empirical baseline: the mean reward across all generations produced by the model for the same prompt. In both methods, rewards are provided by a reward model. Specifically, for each input prompt, the model samples a set of $G$ candidate responses, forming a response group. The reward model assigns a score to each response, and advantages are computed by standardizing rewards within the group: subtracting the group mean and dividing by the group standard deviation. This yields relative, normalized advantage values. As in PPO, GRPO may incorporate a penalty term to prevent the updated policy from drifting too far from the reference policy, typically enforced via a Kullback–Leibler (KL) divergence regularizer to ensure stable updates. A formal definition of the GRPO objective, along with additional details, is provided in App. C.
@@ 4 Proposed Method: Murphy
Figure 2: Overview of Murphy. Given an input prompt ($q_{(\cdot)}$), $G_{1}$ code generations ($o_{(\cdot)}$) are generated and evaluated using a reward function ($r_{(\cdot)}$). Generations that do not achieve the maximum reward are revised based on executor feedback (f), combining the original prompt with the failed output, and re-prompted to generate another $G_{2}$ candidates. This iterative process continues for a fixed number of turns, with rewards from later turns propagated backward. The example illustrates the case with $G_{1}=G_{2}=2$, where $G_{1},G_{2}$ represents the number of rollouts per prompt, per turn, and $\rho(\cdot)$ denotes the credit assignment strategy.
Extending GRPO to multi-turn interaction requires formalizing how environmental feedback informs optimization. We study a feedback-rich code generation setting where model outputs are executed and scored against test suites, yielding two feedback types: (1) quantitative (e.g., proportion of test cases passed) and (2) qualitative (e.g., error traces or failing cases). Murphy extends GRPO by introducing (i) a multi-turn rollout mechanism that conditions generation on feedback, and (ii) a credit-assignment scheme that propagates rewards from future successful turns to earlier attempts.
Multi-turn rollout: For each task prompt, the policy begins at turn 1 by generating $G_{1}$ candidate solutions, each evaluated for reward against the prompt’s test suite. Failed generations are paired with their corresponding feedback and, together with the original prompt and prior outputs, form new conditioning contexts for subsequent turns. At turn $s$, the model generates $G_{s}$ new candidates for each failed case, continuing this feedback–refinement process for a fixed number of turns. This per-prompt, iterative rollout enables progressive improvement conditioned on feedback. Unlike standard GRPO, Murphy delays advantage computation until rollouts in all turns are generated and processed, allowing final rewards to retroactively shape credit assignment across prior turns.
Credit assignment: Once rewards for all turns are computed, Murphy backpropagates rewards from the final turn to earlier turns using a temporal credit-assignment scheme. Although early generations may achieve low initial rewards, incorporating execution feedback in later turns often leads to successful refinements. To ensure that these intermediate steps are properly credited, rewards from later successful generations are distributed to earlier ones according to their contribution to eventual success. After reward redistribution, advantages are computed for each generation, and the Murphy loss is applied to update the policy. Together, these mechanisms extend GRPO to handle multi-turn, feedback-conditioned optimization. Below, we formalize the multi-turn rollout mechanism and credit-assignment process and define the Murphy objective, extending GRPO to incorporate multi-turn feedback.
@@
We denote the current and old policy models by $\pi_{\theta}(\cdot\mid\cdot)$ and $\pi_{\theta_{\text{old}}}(\cdot\mid\cdot)$. $\pi_{\text{ref}}$ denotes the reference model (i.e., the model prior to fine-tuning), which is kept fixed (not updated) throughout training. $\mathcal{P}(Q)$ denotes the distribution over input prompts/questions $Q$, and $O$ the output space. As described earlier, in the first turn, the model generates $G_{1}$ candidate solutions for each prompt. For generations that do not attain the max reward, feedback is obtained from the environment. The model is then re-prompted with the original prompt, its previous output, and the corresponding feedback to produce $G_{s}$ new generations per prompt at turn $s$. This iterative procedure naturally forms a tree structure, where the root corresponds to the original prompt, and each subsequent generation (augmented with feedback) forms a child node, with generations at turn $s+1$ linked to their parent at turn $s$ (see Fig. 2).
@@
Multi-turn rollout formalism.We define a feedback-conditioned rollout tree that captures how generations evolve over $S$ turns. Let $s$ denote the turn index and $G_{s}$ the number of generations per prompt at turn $s$. We use $i_{[1:s]}=(i_{1},\dots,i_{s})$ to index a path in the rollout tree, where $i_{j}$ denotes the branch taken at turn $j$.
Turn 1:The model receives $q_{(1)}\sim\mathcal{P}(Q)$ and generates a response group $\{o_{\{q_{(1)},j\}}\}_{j=1}^{G_{1}}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q_{(1)})$. Each generation is executed to obtain a reward and feedback $(r_{\{q_{(1)},j\}},f_{\{q_{(1)},j\}})$, where $r$ denotes the proportion of tests passed and $f$ contains qualitative executor feedback (e.g., failing tests, error messages).
Turn s+1s{+}1:For any unsolved node at turn $s$ (i.e., one that does not achieve the maximum reward), we form a feedback-conditioned prompt by concatenation: $q_{(s+1,i_{[1:s]})}=[\,q_{(s,i_{[1:s-1]})},\,o_{\{q_{(s,i_{[1:s-1]})},i_{s}\}},\,f_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}\,],$ and sample $G_{s+1}$ children
$\{o_{\{q_{(s+1,i_{[1:s]})},k\}}\}_{k=1}^{G_{s+1}}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q_{(s+1,i_{[1:s]})}),$
each evaluated to obtain $(r_{\{q_{(s+1,i_{[1:s]})},k\}},f_{\{q_{(s+1,i_{[1:s]})},k\}})$. This defines a rollout tree rooted at $q_{(1)}$ (see Fig. 2).
A complete path from the root to a leaf at turn $S$ can be written as
$q_{(1)}\rightarrow o_{\{q_{(1)},i_{1}\}}\rightarrow q_{(2,i_{[1:1]})}\rightarrow\cdots\rightarrow o_{\{q_{(S,i_{[1:S-1]})},i_{S}\}}.$
Leaf nodes at turn $S$ represent the final generations after refinement. We provide a more detailed formal treatment (including explicit Turn 2 expansion and tree construction details) in App. B.
@@
Credit assignment formalism.After all outputs and corresponding rewards are generated and the rollout tree is constructed, we focus on assigning credit from later turns back to earlier ones. To achieve this, we explore two distinct strategies.
@@
Max Reward Strategy (MaRS):The Max Reward Strategy is defined recursively, proceeding from the final turn back to the root. Since the final turn $S$ has no children, the rewards at this turn remain unchanged. We then consider turn s=S1s=S-1. Let $o_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}$ denote a generation at turn s=S1s=S-1 with an associated reward $r_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}$. If this generation already achieves the maximum reward, it has no children; otherwise, its children are defined as:
$C\big(o_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}\big)=\{o_{\{q_{(s+1,i_{[1:s]})},1\}},\ \ldots,\$
$o_{\{q_{(s+1,i_{[1:s]})},G_{S}\}}\}$
@@
$r_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}$ $=\text{max}\Bigg(r_{\{q_{(s,i_{[1:s-1]})},i_{s}\}},\$
$\quad\text{max}\Big(C_{r}(o_{\{q_{(s,i_{[1:s-1]})},i_{s}})\Big)\Bigg)$
@@
This strategy assigns each node the maximum of its own reward and the best reward among its descendants. Intuitively, the descendant maximum represents the best outcome achievable through refinement, while taking the outer maximum ensures a node’s credit never decreases; even when feedback fails to improve performance. This formulation captures the maximum progress achievable from any refinement path starting at that node. The procedure operates recursively in a backward pass: rewards are first updated for all nodes at turn S1S-1 based on their children’s values at turn $S$. This process continues backward through the tree, with each turn $s$ receiving updated rewards based on the values from turn $s+1$, until reaching the root.
Mean Reward Strategy (MeRS):The Mean Reward Strategy follows the same recursive credit assignment structure as the Max Reward Strategy (MaRS), but differs in how rewards are propagated. Inspired by the return computation in REINFORCE (Williams, 1992), MeRS updates each node’s reward by incorporating the discounted mean of its children’s rewards.
For a generation $o_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}$ at turn $s$, let $\mathbb{I}_{\text{unsolved}}$ be an indicator that equals 1 if the problem remains unsolved at this node, and 0 otherwise. The update rule is:
$r_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}$ $=$
$\frac{r_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}+\gamma\cdot\overline{C}_{r}\Big(o_{\{q_{(s,i_{[1:s-1]})},i_{s}\}}\Big)}{\mathbb{I}_{\text{unsolved}}\cdot(S-s)+1}$
where $\gamma\in[0,1]$ is a discount factor controlling the influence of descendant rewards, and $\overline{C}_{r}(\cdot)$ denotes the mean reward over children of unsolved nodes (children of solved nodes are masked out since they represent terminal states).
The denominator equals Ss+1S-s+1 for unsolved nodes and $1$ for solved nodes. This depth-based normalization serves two purposes: (1) it prevents rewards from growing unboundedly during backward propagation, as unsolved nodes can potentially accumulate discounted contributions from up to SsS-s future turns; and (2) it ensures fair comparison between nodes that solve at different turns, a node that solves immediately at turn $s$ retains its full reward, while a node that fails but has successful descendants receives appropriately scaled credit that accounts for the additional refinement steps required. All other aspects of the recursive procedure remain identical to MaRS.
@@
MaRS vs MeRS:MaRS propagates the maximum descendant reward to the earlier turns, emphasizing peak performance, whereas MeRS propagates the mean reward, emphasizing stability and overall consistency. MaRS captures best-case improvement, while MeRS provides a smoother estimate of expected progress. Together, they offer complementary views of feedback-driven credit assignment.
@@
Murphy Objective:Once the rewards are reassigned according to the chosen credit assignment strategy (MaRS or MeRS), advantages are computed similar to standard GRPO. The adjusted rewards serve as the basis for computing normalized advantages at each turn. Conditioned on a prompt $\tilde{q}$ and for $G_{s}$ generations, we normalize each reward by subtracting the mean reward and dividing by the standard deviation of the rewards obtained across the $G_{s}$ generations, yielding the normalized advantage $\hat{A}^{\textsc{Murphy}}_{\tilde{q},i,t}$. Additionally, as defined earlier, each $i$-th generation at turn $s$, $o_{\{q_{(s,i_{[1:s-1]})},i\}}\in O$, corresponds to a complete output trajectory, where $o_{\{q_{(s,i_{[1:s-1]})},i,t\}}$ denotes the $t$-th token and $o_{\{q_{(s,i_{[1:s-1]})},i,<t\}}$ the prefix up to (but excluding) token $t$. We denote the sequence length of the $i$-th generation by $|o_{\{q_{(s,i_{[1:s-1]})},i\}}|$. At each turn, the GRPO objective is applied using these credit-adjusted advantages. This per-turn optimization allows Murphy to incorporate feedback from later turns into earlier updates, effectively extending GRPO to a multi-turn setting. Finally, the divergence between the current and reference policies is captured by $D_{\mathrm{KL}}(\pi_{\theta}\,\|\,\pi_{{\text{ref}}})$, computed over all tokens in the generated sequences. The resulting optimization objective, which integrates credit-assigned rewards, normalized advantages, and KL regularization at each turn, defines the Murphy objective, distinguishing it from the standard GRPO formulation. The full Murphy objective is presented in Sec. 4.
@@
Definition 1. (Murphy Objective) $\mathcal{J}_{\textsc{Murphy}}(\theta)$ $=\mathbb{E}_{q\sim\mathcal{P}(Q)}[\sum_{s,i_{1},\ldots i_{S}}\mathcal{J}_{q_{(s,i_{[1:s]})}}(\theta)]$ Where the per prompt objective at each turn $s$ is: $\mathcal{J}_{q_{(s,i_{[1:s-1]})}}(\theta)$ $=\mathbb{E}_{\{o_{(\tilde{q},i)}\}_{i=1}^{G_{s}}\sim\pi_{\theta_{\text{old}}}(O|\tilde{q})}\Bigg$[ $\sum_{i=1}^{G_{s}}\frac{1}{G_{s}|o_{\{\tilde{q},i\}}|}\sum_{t=1}^{|o_{\{\tilde{q},i\}}|}\Bigg($ $\min\Big(R_{\theta}(\tilde{q},i,t)\hat{A}^{\textsc{Murphy}}_{\tilde{q},i,t},$ $\text{clip}\left(R_{\theta}(\tilde{q},i,t),\,1-\epsilon,\,1+\epsilon\right)\cdot$ $\hat{A}^{\textsc{Murphy}}_{\tilde{q},i,t}\Big)-\beta D_{\mathrm{KL}}(\pi_{\theta}\,\|\,\pi_{{\text{ref}}})\Bigg)\Bigg]$ $\text{with,~}\tilde{q}$ $=q_{(s,i_{[1:s-1]})}$ $R_{\theta}(\tilde{q},i,t)$ $=\frac{\pi_{\theta}(o_{\{\tilde{q},i,t\}}\mid\tilde{q},o_{\{\tilde{q},i,<t\}})}{\pi_{\theta_{\text{old}}}(o_{\{\tilde{q},i,t\}}\mid\tilde{q},o_{\{\tilde{q},i,<t\}})}$ And, $\hat{A}^{\textsc{Murphy}}_{\tilde{q},i,t}$ denotes the advantage.